OpenStudy (anonymous):

What is the solution to the following system of equations? 2x + 3y = 6 x + y = 4 The solution of the system of equations is

OpenStudy (haseeb96):

step # 1 call the equations as 2x+3y =6.............. eq # 1 x+y=4.............eq #2 step #2 isolate the x from the most simplest equation I took equation 2 x = 4-y ....... eq#3 I called this equation as equation 3 Step # 3 now put the x =4-y into the equation 1 at the place of x so equation 1 becomes 2 (4-y) +3y = 6 8-2y + 3y = 6 8 + y =6 y=6-8 y=-2 now u will put y=-2 into equation 3 at the place of y since x = 4-y x = 4 - (-2) x = 4+2 x= 6 so x= 6 and y =-2 this is your solution

OpenStudy (haseeb96):

@mrkdc2 if you have any question regarding this explanation ask me.

OpenStudy (anonymous):

no i dont thank you

OpenStudy (haseeb96):

Welcome